OpenStudy (dgraves11):

A 20-foot wire connects the top of a utility pole to a point on the ground located 10 feet away from the base of the pole.

OpenStudy (legomyego180):



OpenStudy (legomyego180):

Height of the pole = \(\sqrt{20^2-10^2}\) by the pythagorean theorum.

OpenStudy (arjunsharma1065):

@legomyego180 Doesn't \[a ^{2} + b^{2} = c^{2}\]? So shouldn't the height of the pole really be \[\sqrt{a^{2} + b^{2}}\]

OpenStudy (legomyego180):

 \[a^2+b^2=c^2\] Rewriting this to solve for b: \[a^2+b^2=c^2\] \[b^2=c^2-a^2\] \[b=\pm \sqrt{c^2-a^2}\]

OpenStudy (arjunsharma1065):

@legomyego180 shoot ur right, sorry i was braindead